Let T(x,y)=(3x−8y,−4x−9y)T(x,y)=(3x−8y,−4x−9y). Find x such that T(x)=(−9,−2)T(x→)=(−9,−2).
Answer #1

Given that TOY) = (34-84, 41-gy) We have to find (ory) such that T(my) = (-9,-2) SO fi 32-8y = -9 -40-gy = -2 Oxs + x3 ! 28 -If you don't understand please comment